# MoDora: Tree-Based Semi-Structured Document Analysis System

### v1

@@ 1. Introduction
Figure 1. Example Semi-Structured Document Analysis.
@@ 2.1. Semi-Structured Document
2.2. Semi-Structured Document Analysis
@@ 7. Micro Benchmark Preparation
Document Diversity Metrics. Since basic metrics (e.g., number of pages and elements) fail to capture a document’s layout complexity and element diversity, we introduce two new metrics to characterize structural richness: Element Diversity ($ED$), which measures the diversity of element types, and Layout Diversity ($LD$), which quantifies the variance in the components’ locational distribution.
(1) Element Diversity (ED). A document dominated by any single type (e.g., tables, text) tends to exhibit low diversity. Thus we define Element Diversity ($ED$) as the normalized standard variance in the counts of the five element categories (see Section 2), i.e.,
@@
where $t,b,c,i,s$ denote text, table, chart, image, and supplement respectively and $n$ is the number of total elements. A larger $ED$ value indicates lower element diversity.
(2) Layout Diversity (LD). While $ED$ measures content variety, it does not reflect structural composition. We therefore define Layout Diversity ($LD$) to quantify the average number of independent components per page, i.e., $LD=\frac{N_{\text{comp}}}{N_{\text{page}}}$, where $N_{\text{comp}}$ denotes the number of independent components and $N_{\text{page}}$ the total number of pages. An independent component refers to a contiguous sequence of elements of the same type; for example, four adjacent tables are counted as one independent component. A higher $LD$ indicates richer layout organization and stronger structural variation.
Dataset Preparation. To construct the MMDA benchmark, we begin by sampling 200k candidate documents from over one million real-world documents, spanning academic papers, magazines, financial reports, and presentation slides. Benchmark data is selected from these candidates in three steps. First, to emphasize the layout structures, we annotate each document by overlaying translucent rectangles of maximally distinguishable colors onto regions corresponding to different element types. Second, we compute embeddings using ViT-Base-Patch16-224 (Wu et al., 2020) by the visual features, and apply K-means clustering (with 1,000 clusters, two documents per cluster) to identify 2,000 layout-diverse ones. Third, we rank these documents by their Element Diversity ($ED$) scores and retain the top 537 after filtering out low-quality samples (e.g., documents with blurred or ambiguous content).
@@ 8.3. Fine-Grained Result Analysis
Comparisons using ZenDB benchmarks. We further evaluate MoDora against ZenDB (24) using the same datasets and evaluation metrics reported in (24). As shown in Table 4, MoDora consistently delivers substantially stronger performance, demonstrating robust generalizability across seven benchmarks (Civic, Qasper, Contract, DUDE, MMDA, M3DocVQA, and MP-DocVQA) and a diverse set of evaluation metrics (precision, recall, f1, ACNLS, AIC-CC).

### v2

# arXiv:2602.23061v2[cs.IR] 27 Feb 2026

@@ 1. Introduction
Figure 1. Example Semi-Structured Document Analysis.
@@ 2.1. Semi-Structured Document
$2.$ Semi-Structured Document Analysis
@@ 7. Micro Benchmark Preparation
Document Diversity Metrics. Since basic metrics (e.g., number of pages and elements) fail to capture a document’s layout complexity and element diversity, we introduce two new metrics to characterize structural richness: Element Diversity (EDED), which measures the diversity of element types, and Layout Diversity (LDLD), which quantifies the variance in the components’ locational distribution.
(1) Element Diversity (ED). A document dominated by any single type (e.g., tables, text) tends to exhibit low diversity. Thus we define Element Diversity (EDED) as the normalized standard variance in the counts of the five element categories (see Section 2), i.e.,
@@
where $t,b,c,i,s$ denote text, table, chart, image, and supplement respectively and $n$ is the number of total elements. A larger EDED value indicates lower element diversity.
(2) Layout Diversity (LD). While EDED measures content variety, it does not reflect structural composition. We therefore define Layout Diversity (LDLD) to quantify the average number of independent components per page, i.e., $LD=\frac{N_{\text{comp}}}{N_{\text{page}}}$, where $N_{\text{comp}}$ denotes the number of independent components and $N_{\text{page}}$ the total number of pages. An independent component refers to a contiguous sequence of elements of the same type; for example, four adjacent tables are counted as one independent component. A higher LDLD indicates richer layout organization and stronger structural variation.
Dataset Preparation. To construct the MMDA benchmark, we begin by sampling 200k candidate documents from over one million real-world documents, spanning academic papers, magazines, financial reports, and presentation slides. Benchmark data is selected from these candidates in three steps. First, to emphasize the layout structures, we annotate each document by overlaying translucent rectangles of maximally distinguishable colors onto regions corresponding to different element types. Second, we compute embeddings using ViT-Base-Patch16-224 (Wu et al., 2020) by the visual features, and apply K-means clustering (with 1,000 clusters, two documents per cluster) to identify 2,000 layout-diverse ones. Third, we rank these documents by their Element Diversity (EDED) scores and retain the top 537 after filtering out low-quality samples (e.g., documents with blurred or ambiguous content).
@@ 8.3. Fine-Grained Result Analysis
Comparisons using ZenDB benchmarks. We further evaluate MoDora against ZenDB (24) using the same datasets and evaluation metrics reported in (24). As shown in Table 4, MoDora consistently delivers substantially stronger performance, demonstrating robust generalizability across seven benchmarks (Civic, Qasper, Contract, DUDE, MMDA, M3DocVQA, and MP-DocVQA) and a diverse set of evaluation metrics (precision, recall, f1, ACNLS, AIC-$C$).